Suppose that marginal income tax rates are as follows:
 

Income levelMarginal tax rate
$0–$10,00010%
$10,000–$50,00020%
Greater than $50,00035%

 


Also suppose that the taxes for Social Security and Medicare together are 7.65%, and that they are applied to income up to $100,000.


If your income is $123,000, what is your total income tax bill and how much will you have paid in Social Security and Medicare taxes?

Total income tax:   $   
Social Security and Medicare taxes:   $ 
Answer #1

Suppose that there is a person whose income is \(\$ 123,000\) The taxes for social security and Medicare together are \(7.65 \%\) and they are applied up to the income of \(\$ 100,000\).

Find the total income tax.

total income \(\operatorname{tax}=10 \%\) of \((10,000-0)+20 \%\) of \((50,000-10,000)+\)

$$ \begin{array}{l} 35 \% \text { of }(123,000-50,000) \\ =10 \% \times 10,000+20 \% \times 40,000+35 \% \times(73,000) \\ =1,000+8,000+25,550 \\ =\$ 34,550 \end{array} $$

Thus, the total income \(\operatorname{tax}\) is \(\$ 34,550\)


The social security and Medicare tax is calculated as follows:

\(=7.65 \%(\$ 100,000)\)

\(=\$ 7650\)

The security and medicare taxes is \(\$ 7650\)


answered by: Serpent